# A compact QUBO encoding of computational logic formulae demonstrated on cryptography constructions

### v1

@@ Abstract
We aim to advance the state-of-the-art in Quadratic Unconstrained Binary Optimization formulation with a focus on cryptography algorithms. As the minimal QUBO encoding of the linear constraints of optimization problems emerges as the solution of integer linear programming (ILP) problems, by solving special boolean logic formulas (like ANF and DNF) for their integer coefficients it is straightforward to handle any normal form, or any substitution for multi-input AND, OR or XOR operations in a QUBO form. To showcase the efficiency of the proposed approach we considered the most widespread cryptography algorithms including AES-128/192/256, MD5, SHA1 and SHA256. For each of these, we achieved QUBO instances reduced by thousands of logical variables compared to previously published results, while keeping the QUBO matrix sparse and the magnitude of the coefficients low. In the particular case of AES-256 cryptography function we obtained more than 8×8$\times$8 × reduction in variable count compared to previous results. The demonstrated reduction in QUBO sizes notably increases the vulnerability of cryptography algorithms against future quantum annealers, capable of embedding around $30$ thousands of logical variables.
@@ 1 Introduction
As SAT–to–QUBO encoding strategies turned out to have a significant impact on the solution quality of quantum annealing [47, 48], numerous research work have been conducting on QUBO transformation methods [49, 50, 28, 51, 24, 30, 52, 53, 54, 55, 31]. According to the study presented in Ref. [28] the SAT–to–QUBO transform can be formulated as a problem in Satisfiability or Optimization Modulo Theories (SMT/OMT) [56, 57] on the theory of linear rational arithmetic. By encoding the whole SAT problem in one step would typically require the introduction of many additional ancillary boolean variables. These extra variables would result in very large SMT/OMT formulas: solving the resulting problem would become possibly even harder than the original SAT problem. In order to overcome this issue scalable “divide-and-conquer” approach was proposed for the SAT–to–QUBO transformation. In this regard, the special forms of SAT instances mentioned above can be used to engineer encoding strategies to reduce the variable count in the formulation of CO problems [58]. The work of Ref. [27], for example, presented two novel QUBO formulations for $k$-SAT and Hamiltonian Cycles that scale significantly better than preceding approaches. For $k$-SAT the growth of the QUBO matrix was reduced from $\mathcal{O}(k)$ to $\mathcal{O}(\textrm{log}(k))$, while for Hamiltonian Cycles the new approach gives a QUBO matrix scaling linearly in the number of edges and logarithmically in the number of nodes. The reported improvement was achieved by an iterative squaring technique for the CNF form of SAT. Additionally, a systematic meta-algorithm was introduced in the study of Ref. [59], facilitating the design of automatized QUBO encoding of $3$-SAT problems. The authors of this work pointed out the importance of patterns in SAT formulations and use their properties in algorithmic methods to encode QUBO instances. The study in Ref. [60] introduced a logarithmicaly scaling method to encode equality and inequality constraints that was further enhanced by research of Ref. [61] providing an asymptotic improvement in the number of variables.
@@ 2 The theory of finding QUBO Encoding Patterns
A QUBO problem with $N$ binary variables is represented by a minimization (or maximization) of
@@
In principle, the valid values of the variables $\mathbf{s}$ are determined via a problem specific binary-input function $\mathbf{s}_{\textrm{valid}}=\mathbf{s}_{\textrm{valid}}(\mathbf{x})$. The related literature provides countless examples to construct the function $\mathbf{s}_{\textrm{valid}}(\mathbf{x})$ [62, 63, 64]. Usually these examples involve the reduction of higher order terms, such as the transform of a 3-SAT term $x_{1}\land x_{2}\land x_{3}$ into QUBO by introducing a substitution variable $s=s_{\textrm{valid}}(\mathbf{x})=x_{1}\cdot x_{2}$ and append a penalty term to the QUBO formula giving a zero energy only for $s=s_{\textrm{valid}}(\mathbf{x})$ and is greater than zero otherwise. For this specific case Rosenberg showed that $g(\mathbf{x},s)=3s+x_{1}x_{2}-2x_{1}s-2x_{2}s$ is a well chosen function to fulfill these requirements [65]. However, for larger problems the outlined strategy to define substitution variables turns to be wasteful in the variable count as one can find better encoding by considering larger portions of the SAT formula and finding nontrivial substitutions and penalty terms as it was demonstarted in the work of Ref. [66]. Here we describe our general strategy of finding compact encoding QUBO formulae. To this end we introduce a quadratic formula $g(\mathbf{x},\mathbf{s})$ as a function of binary substitution variables encompassed by the vector $\mathbf{s}$. We expect the quadratic function $g(\mathbf{x},\mathbf{s})$ to take on a return value 00 for any min-term input $\mathbf{x}_{\textrm{min}}$ and for the substitution variables $\mathbf{s}_{\textrm{valid}}(\mathbf{x}_{\textrm{min}})$:
@@
For any other binary inputs $g(\mathbf{x},\mathbf{s})$ should give values larger than 00:
@@
where $\mathbf{s}_{i}$ is the $i$-th bit-string of the substitution variables during the exhaustive search, such that $\mathbf{s}_{i}$ is nothing else than the binary representation of the integer $i$. Then, the ILP problem can be formulated as follows
@@
where we have replaced the polynomial $g$ with $\tilde{g}$ and used $\wedge$ clauses instead of $\vee$ operators. This transformation is justified as valid substitution variable maps are determined by clauses being satisfied with $k_{i}=0$. The conjunctive clauses in (10) and (11) are straightforward to convert into ILP as all of the involved clauses need to be satisfied simultaneously, resulting in a set of linear constraints of the form of Eq. (3) with respect to the unknown coefficients in the polynomial $g$. Regarding the constructed ILP, the condition (7) ensures two important properties. Firstly, as the slack variables $t_{i}$ are greater than or equal to $1$ and $k_{i}$ is either 00 or $1$, constraint (5) will be satisfied when the constructed ILP is solved. Secondly, by summing up the values of $k_{i}$ in (11) we can deduce the number of bit-strings $\mathbf{s}_{i}$ satisfying Eq. (4). If $\sum_{i}k_{i}=2^{N_{s}}$ holds on, then the ILP has not found any valid substitution variable vector satisfying (4), and we need to increase $N_{s}$. When $\sum_{i}k_{i}=2^{N_{s}}-1$, then there is only a single bit-string $\mathbf{s}$ satisfying (4) and the number of the substitution variables should be considered optimal. Finally, we can assume that the number of the substitution variables can be decreased then $\sum_{i}k_{i}<2^{N_{s}}-1$. The equation $\sum_{i}k_{i}=2^{N_{s}}-1$ implicitly determines the optimal number $N_{s}$ of the substitution variables. The solution of the ILP problem implicitly also determines the unique substitution variable vector $\mathbf{s}=\mathbf{s}_{\textrm{valid}}(\mathbf{x}_{\textrm{min}})=\mathbf{s}_{i}$ corresponding to $k_{i}=0$ in Eq (7).
@@
$\begin{split}\mathcal{P}_{\textrm{multi-min-terms}}=\\ \bigg{(}\tilde{g}(\mathbf{x}_{\textrm{min}}^{(1)},\mathbf{s}_{0})=0\quad\wedge\quad&\dots\quad\wedge\quad\tilde{g}(\mathbf{x}_{\textrm{min}}^{(1)},\mathbf{s}_{2^{N_{s}}-1})=0\bigg{)}\\ &\wedge\\ \bigg{(}\tilde{g}(\mathbf{x}_{\textrm{min}}^{(2)},\mathbf{s}_{0})=0\quad\wedge\quad&\dots\quad\wedge\quad\tilde{g}(\mathbf{x}_{\textrm{min}}^{(2)},\mathbf{s}_{2^{N_{s}}-1})=0\bigg{)}\\ &\wedge\\ \dots{}\end{split}$ (12)
@@ 3.1 Parity encoding: multi-input XOR clauses
Let us first examine parity encoding. The very first thing to decide is the way the substitution variables are represented. The idea here is that the parity function decides on a sum of input variables whether it is odd (return value $1$) or not (return value 00). To turn this function into a QUBO form (supposing we limit the input variables to binary ones) we search for the minimum of
@@
where the substitution variable $T$ can take on values from the set of odd integers greater than 00. If the sum of the variables is odd, one can find a suitable value for $T$ to cancel the sum ending in a minimal value of 00 for the quadratic expression, indicating that the sum is odd. Otherwise, the substitution variable $T$ can not cancel the sum, giving nonzero output for the quadratic expression even at its minimum. We now look for the optimal binary representation of $T$. Odd integers can be given in a form $T=2t-1$ with $1\leq t\leq t_{max}$. $t_{max}$ is determined according to the maximal value of $\sum_{i}x_{i}$ that $T$ needs to cancel. In case we have $N$ binary variables in the sum, we get $t_{max}=\lceil N/2\rceil$. Then, if $t_{\max}$ is equal to a power of $2$, the most optimal encoding of $t$ is simply taking its binary expansion:
@@
with a number of substitution variables $N_{subs}=\log_{2}(t_{max})$. In a more general case when $t_{max}$ is not a power of $2$ we propose a modified expression to encode a slack variable covering the needed range:
@@
where $N_{subs}=\lceil\log_{2}(t_{max})\rceil$. In this expression the most significant binary variable $s_{N_{subs}-1}$ has a coefficient whose value is relaxed compared to a conventional binary number representation in front of $s_{N_{subs}-1}$. The relaxed binary expansion is general, can be used either when $t_{max}$ is a power of $2$ or not. With this modified binary expansion form, however, there occurs no overflows in the covered range determined by $[1,t_{\max}]$.
@@ 3.2 Range constraints: multi-input OR clauses
where a suitable substitution variable $1\leq T\leq N$ can be found for the equality. In case having such $T$, the multi-input OR function (17) can be satisfied to give True output. Otherwise, Eq.(17) is giving output False and (18) gets violated. As the binary and integer variables can change only by $1$, (16) can take on only non-negative values, while giving 00 only when fulfilling (18) with $T=2t-1$ or $T=2t-2$. Therefore, the substitution variable $t$ is required to span a shorter range of $1\leq t\leq t_{max}$, where $t_{max}=\lceil(N+1)/2\rceil$. The shorter range requires one less bit in the binary expansion of $t$ resulting in saving one bit (i.e. one QUBO variable) in the encoding when using (16) instead of 13). For the binary expansion of $t$ we can use (15) to cover the range $1\leq t\leq t_{max}$. An important point here is to understand the mechanism of (16) on saving one bit in the encoding of the substitution variable. As one can see, a single combination of substitution variables in this formula is used to encode two different values at once, which are $12t-$ 1 and $22t-$ 2 ($t$ is expressed in terms of substitution variables $s_{i}$ via a reduced binary expansion of Eq. (15)). This dualism is the key reason for the saving of one bit in the encoding. In the case of even $N$ a subtle detail to notice is that $12t_{max}-1=N+$ 1, seemingly overflowing the range limit $N$. However, the left side of Eq. (16) will becomes zero for $t=t_{max}$, thus the QUBO encoding (16) will not take on negative values or any false minima.
@@
where functions $f_{1}(\{s_{i}\})$ and $f_{2}(\{s_{i}\})$ must cover all elements in the set $\{1,2,3,\dots,N\}$. When $t_{max}$ is not a power of $2$, the encoding of the combinations of the binary substitution variables $\{s_{i}\}$ is larger than the range needed to be spanned. Therefore, $f_{1}$ and $f_{2}$ don’t need to give a unique cover of the spanned range, they might coincide for some inputs.
Equation (19) is the most general form to encode the multi-input OR formula. The terms on the right and left must differ precisely by $1$ (or equal to each other), otherwise the product might give negative values for some inputs $\{x_{i}\}$ obscuring the QUBO encoding by providing incorrect minima. A simple example for $f_{1}(\{s_{i}\})$ and $f_{2}(\{s_{i}\})$ can be given for the case of a multi-input OR formula with $N=10$ variables $x_{i}$. In this case $t_{max}=6$, thus $N_{s}=3$ and
@@
Range constraints expressed with strict inequalities (i.e. $<$ and $>$) can be transformed into the form (22) by subtracting $1$ from $H$ or adding $1$ to $L$. In this case the range $T$ to be spanned by substitution variables is determined by $T=H-L+1$ (the lowest and highest values are also included in the range). As the population count $\sum x_{i}$ can exceed $H$ or can be less than $L$, extra caution needs to be payed preventing from overstepping the boundaries $L$ and $H$ by the minima of the QUBO encoded form of the range constraint, otherwise it will convey false minima. For a range $T$ (being either even or odd) the function $f_{1}$ can be encoded as
@@
with $N_{subs}=\lfloor\log_{2}(T)\rfloor$ and similarly to Eq. (15), a relaxed coefficient in front of the most significant bit $s_{N_{subs}-1}$. The distance between any pair of nearest values of $f_{1}$ – ensured by the least significant bits – never exceeds $2$. Some $\{s_{i}\}$ combinations can coincide in the resulting value of $f_{1}$ (for even $T$) or differ only by $1$ (for odd $T$). Therefore, a suitable $f_{2}$ can be constructed by simply offsetting $f_{1}$ by one:
@@
Functions $f_{1}$ and $f_{2}$ cover the whole range of $\{L,L+1,\dots,H\}$ with alternating values coming from $f_{1}$ and $f_{2}$. An important detail here is that when $f_{1}=H$ then $f_{2}$ will take on the value of $H+1$, overstepping the given range. However, similarly to the QUBO encoding (16), the left side of Eq. (19) will prevent from picking up negative values (or false minima) by the QUBO objective function. Further valid $f_{1}$ and $f_{2}$ encodings can be generated when the relaxation of the coefficients is applied to the other coefficients as well, provided that the difference between any pair of nearest elements among the possible values of $f_{1}$ will not exceed $2$. Because of this reason, the relaxation of the least significant bit is not allowed. The rule of thumb of the relaxation is that the coefficient in front of any $s_{i}$ must be less or equal to the sum of the coefficients in front of $s_{j}$-s with $1\leq j<i$ or larger than this sum at most by $2$, provided that the coefficients are in ascending order from $i=0$ to $i=N_{subs}-1$. The amount of the relaxation, i.e. the accumulated decrease of the coefficients, equals to $2^{N_{subs}+1}-T$, as $2^{N_{subs}+1}$ is the maximum of the range that can be spanned with $N_{subs}$ binary variables. Following this rule and Eq. (24) one can construct a plenty of valid encodings $f_{1}$ and $f_{2}$.

### v2

@@ Abstract
We aim to advance the state-of-the-art in Quadratic Unconstrained Binary Optimization formulation with a focus on cryptography algorithms. As the minimal QUBO encoding of the linear constraints of optimization problems emerges as the solution of integer linear programming (ILP) problems, by solving special boolean logic formulas (like ANF and DNF) for their integer coefficients it is straightforward to handle any normal form, or any substitution for multi-input AND, OR or XOR operations in a QUBO form. To showcase the efficiency of the proposed approach we considered the most widespread cryptography algorithms including AES-128/192/256, MD5, SHA1 and SHA256. For each of these, we achieved QUBO instances reduced by thousands of logical variables compared to previously published results, while keeping the QUBO matrix sparse and the magnitude of the coefficients low. In the particular case of AES-256 cryptography function we obtained more than $8\times$ reduction in variable count compared to previous results. The demonstrated reduction in QUBO sizes notably increases the vulnerability of cryptography algorithms against future quantum annealers, capable of embedding around $30$ thousands of logical variables.
@@ 1 Introduction
As SAT–to–QUBO encoding strategies turned out to have a significant impact on the solution quality of quantum annealing [47, 48], numerous research work have been conducting on QUBO transformation methods [49, 50, 28, 51, 24, 30, 52, 53, 54, 55, 31]. According to the study presented in Ref. [28] the SAT–to–QUBO transform can be formulated as a problem in Satisfiability or Optimization Modulo Theories (SMT/OMT) [56, 57] on the theory of linear rational arithmetic. By encoding the whole SAT problem in one step would typically require the introduction of many additional ancillary boolean variables. These extra variables would result in very large SMT/OMT formulas: solving the resulting problem would become possibly even harder than the original SAT problem. In order to overcome this issue scalable “divide-and-conquer” approach was proposed for the SAT–to–QUBO transformation. In this regard, the special forms of SAT instances mentioned above can be used to engineer encoding strategies to reduce the variable count in the formulation of CO problems [58]. The work of Ref. [27], for example, presented two novel QUBO formulations for $k$-SAT and Hamiltonian Cycles that scale significantly better than preceding approaches. For $k$-SAT the growth of the QUBO matrix was reduced from $\mathcal{O}(k)$ to $\mathcal{O}(\textrm{log}(k))$, while for Hamiltonian Cycles the new approach gives a QUBO matrix scaling linearly in the number of edges and logarithmically in the number of nodes. The reported improvement was achieved by an iterative squaring technique for the CNF form of SAT. Additionally, a systematic meta-algorithm was introduced in the study of Ref. [59], facilitating the design of automatized QUBO encoding of 33-SAT problems. The authors of this work pointed out the importance of patterns in SAT formulations and use their properties in algorithmic methods to encode QUBO instances. The study in Ref. [60] introduced a logarithmicaly scaling method to encode equality and inequality constraints that was further enhanced by research of Ref. [61] providing an asymptotic improvement in the number of variables.
@@ 2 The theory of finding QUBO Encoding Patterns
A QUBO problem with NN binary variables is represented by a minimization (or maximization) of
@@
In principle, the valid values of the variables $\mathbf{s}$ are determined via a problem specific binary-input function $\mathbf{s}_{\textrm{valid}}=\mathbf{s}_{\textrm{valid}}(\mathbf{x})$. The related literature provides countless examples to construct the function $\mathbf{s}_{\textrm{valid}}(\mathbf{x})$ [62, 63, 64]. Usually these examples involve the reduction of higher order terms, such as the transform of a 3-SAT term $x_{1}\land x_{2}\land x_{3}$ into QUBO by introducing a substitution variable $s=s_{\textrm{valid}}(\mathbf{x})=x_{1}\cdot x_{2}$ and append a penalty term to the QUBO formula giving a zero energy only for $s=s_{\textrm{valid}}(\mathbf{x})$ and is greater than zero otherwise. For this specific case Rosenberg showed that $g(\mathbf{x},s)=3s+x_{1}x_{2}-2x_{1}s-2x_{2}s$ is a well chosen function to fulfill these requirements [65]. However, for larger problems the outlined strategy to define substitution variables turns to be wasteful in the variable count as one can find better encoding by considering larger portions of the SAT formula and finding nontrivial substitutions and penalty terms as it was demonstarted in the work of Ref. [66]. Here we describe our general strategy of finding compact encoding QUBO formulae. To this end we introduce a quadratic formula $g(\mathbf{x},\mathbf{s})$ as a function of binary substitution variables encompassed by the vector $\mathbf{s}$. We expect the quadratic function $g(\mathbf{x},\mathbf{s})$ to take on a return value 0 for any min-term input $\mathbf{x}_{\textrm{min}}$ and for the substitution variables $\mathbf{s}_{\textrm{valid}}(\mathbf{x}_{\textrm{min}})$:
@@
For any other binary inputs $g(\mathbf{x},\mathbf{s})$ should give values larger than 0:
@@
where $\mathbf{s}_{i}$ is the $i$-th bit-string of the substitution variables during the exhaustive search, such that $\mathbf{s}_{i}$ is nothing else than the binary representation of the integer ii. Then, the ILP problem can be formulated as follows
@@
where we have replaced the polynomial $g$ with $\tilde{g}$ and used $\wedge$ clauses instead of $\vee$ operators. This transformation is justified as valid substitution variable maps are determined by clauses being satisfied with $k_{i}=0$. The conjunctive clauses in (10) and (11) are straightforward to convert into ILP as all of the involved clauses need to be satisfied simultaneously, resulting in a set of linear constraints of the form of Eq. (3) with respect to the unknown coefficients in the polynomial $g$. Regarding the constructed ILP, the condition (7) ensures two important properties. Firstly, as the slack variables $t_{i}$ are greater than or equal to 11 and $k_{i}$ is either 0 or 11, constraint (5) will be satisfied when the constructed ILP is solved. Secondly, by summing up the values of $k_{i}$ in (11) we can deduce the number of bit-strings $\mathbf{s}_{i}$ satisfying Eq. (4). If $\sum_{i}k_{i}=2^{N_{s}}$ holds on, then the ILP has not found any valid substitution variable vector satisfying (4), and we need to increase $N_{s}$. When $\sum_{i}k_{i}=2^{N_{s}}-1$, then there is only a single bit-string $\mathbf{s}$ satisfying (4) and the number of the substitution variables should be considered optimal. Finally, we can assume that the number of the substitution variables can be decreased then $\sum_{i}k_{i}<2^{N_{s}}-1$. The equation $\sum_{i}k_{i}=2^{N_{s}}-1$ implicitly determines the optimal number $N_{s}$ of the substitution variables. The solution of the ILP problem implicitly also determines the unique substitution variable vector $\mathbf{s}=\mathbf{s}_{\textrm{valid}}(\mathbf{x}_{\textrm{min}})=\mathbf{s}_{i}$ corresponding to $k_{i}=0$ in Eq (7).
@@
[…]
@@ 3.1 Parity encoding: multi-input XOR clauses
Let us first examine parity encoding. The very first thing to decide is the way the substitution variables are represented. The idea here is that the parity function decides on a sum of input variables whether it is odd (return value 11) or not (return value 0). To turn this function into a QUBO form (supposing we limit the input variables to binary ones) we search for the minimum of
@@
where the substitution variable $T$ can take on values from the set of odd integers greater than 0. If the sum of the variables is odd, one can find a suitable value for $T$ to cancel the sum ending in a minimal value of 0 for the quadratic expression, indicating that the sum is odd. Otherwise, the substitution variable $T$ can not cancel the sum, giving nonzero output for the quadratic expression even at its minimum. We now look for the optimal binary representation of $T$. Odd integers can be given in a form T=2t1T=2t-1 with $1\leq t\leq t_{max}$. $t_{max}$ is determined according to the maximal value of $\sum_{i}x_{i}$ that $T$ needs to cancel. In case we have NN binary variables in the sum, we get $t_{max}=\lceil N/2\rceil$. Then, if $t_{\max}$ is equal to a power of 22, the most optimal encoding of $t$ is simply taking its binary expansion:
@@
with a number of substitution variables $N_{subs}=\log_{2}(t_{max})$. In a more general case when $t_{max}$ is not a power of 22 we propose a modified expression to encode a slack variable covering the needed range:
@@
where $N_{subs}=\lceil\log_{2}(t_{max})\rceil$. In this expression the most significant binary variable $s_{N_{subs}-1}$ has a coefficient whose value is relaxed compared to a conventional binary number representation in front of $s_{N_{subs}-1}$. The relaxed binary expansion is general, can be used either when $t_{max}$ is a power of 22 or not. With this modified binary expansion form, however, there occurs no overflows in the covered range determined by $1,t_{\max}]$.
@@ 3.2 Range constraints: multi-input OR clauses
where a suitable substitution variable $1\leq T\leq N$ can be found for the equality. In case having such $T$, the multi-input OR function (17) can be satisfied to give True output. Otherwise, Eq.(17) is giving output False and (18) gets violated. As the binary and integer variables can change only by 11, (16) can take on only non-negative values, while giving 0 only when fulfilling (18) with T=2t1T=2t-1 or T=2t2T=2t-2. Therefore, the substitution variable $t$ is required to span a shorter range of $1\leq t\leq t_{max}$, where $t_{max}=\lceil(N+1)/2\rceil$. The shorter range requires one less bit in the binary expansion of $t$ resulting in saving one bit (i.e. one QUBO variable) in the encoding when using (16) instead of 13). For the binary expansion of $t$ we can use (15) to cover the range $1\leq t\leq t_{max}$. An important point here is to understand the mechanism of (16) on saving one bit in the encoding of the substitution variable. As one can see, a single combination of substitution variables in this formula is used to encode two different values at once, which are 2t12t-1 and 2t22t-2 ($t$ is expressed in terms of substitution variables $s_{i}$ via a reduced binary expansion of Eq. (15)). This dualism is the key reason for the saving of one bit in the encoding. In the case of even NN a subtle detail to notice is that $2t_{max}-1=N+1$, seemingly overflowing the range limit NN. However, the left side of Eq. (16) will becomes zero for $t=t_{max}$, thus the QUBO encoding (16) will not take on negative values or any false minima.
@@
where functions $f_{1}(\{s_{i}\})$ and $f_{2}(\{s_{i}\})$ must cover all elements in the set $\{1,2,3,\dots,N\}$. When $t_{max}$ is not a power of 22, the encoding of the combinations of the binary substitution variables $\{s_{i}\}$ is larger than the range needed to be spanned. Therefore, $f_{1}$ and $f_{2}$ don’t need to give a unique cover of the spanned range, they might coincide for some inputs.
Equation (19) is the most general form to encode the multi-input OR formula. The terms on the right and left must differ precisely by 11 (or equal to each other), otherwise the product might give negative values for some inputs $\{x_{i}\}$ obscuring the QUBO encoding by providing incorrect minima. A simple example for $f_{1}(\{s_{i}\})$ and $f_{2}(\{s_{i}\})$ can be given for the case of a multi-input OR formula with $N=10$ variables $x_{i}$. In this case $t_{max}=6$, thus $N_{s}=3$ and
@@
Range constraints expressed with strict inequalities (i.e. << and >>) can be transformed into the form (22) by subtracting 11 from $H$ or adding 11 to $L$. In this case the range $T$ to be spanned by substitution variables is determined by T=HL+1T=H-L+1 (the lowest and highest values are also included in the range). As the population count $\sum x_{i}$ can exceed $H$ or can be less than $L$, extra caution needs to be payed preventing from overstepping the boundaries $L$ and $H$ by the minima of the QUBO encoded form of the range constraint, otherwise it will convey false minima. For a range $T$ (being either even or odd) the function $f_{1}$ can be encoded as
@@
with $N_{subs}=\lfloor\log_{2}(T)\rfloor$ and similarly to Eq. (15), a relaxed coefficient in front of the most significant bit $s_{N_{subs}-1}$. The distance between any pair of nearest values of $f_{1}$ – ensured by the least significant bits – never exceeds 22. Some $\{s_{i}\}$ combinations can coincide in the resulting value of $f_{1}$ (for even $T$) or differ only by 11 (for odd $T$). Therefore, a suitable $f_{2}$ can be constructed by simply offsetting $f_{1}$ by one:
@@
Functions $f_{1}$ and $f_{2}$ cover the whole range of $\{L,L+1,\dots,H\}$ with alternating values coming from $f_{1}$ and $f_{2}$. An important detail here is that when $f_{1}=H$ then $f_{2}$ will take on the value of $H+1$, overstepping the given range. However, similarly to the QUBO encoding (16), the left side of Eq. (19) will prevent from picking up negative values (or false minima) by the QUBO objective function. Further valid $f_{1}$ and $f_{2}$ encodings can be generated when the relaxation of the coefficients is applied to the other coefficients as well, provided that the difference between any pair of nearest elements among the possible values of $f_{1}$ will not exceed 22. Because of this reason, the relaxation of the least significant bit is not allowed. The rule of thumb of the relaxation is that the coefficient in front of any $s_{i}$ must be less or equal to the sum of the coefficients in front of $s_{j}$-s with $1\leq j<i$ or larger than this sum at most by 22, provided that the coefficients are in ascending order from $i=0$ to $i=N_{subs}-1$. The amount of the relaxation, i.e. the accumulated decrease of the coefficients, equals to $2^{N_{subs}+1}-T$, as $2^{N_{subs}+1}$ is the maximum of the range that can be spanned with $N_{subs}$ binary variables. Following this rule and Eq. (24) one can construct a plenty of valid encodings $f_{1}$ and $f_{2}$.